# Factorisation conditions and causality for local measurements in QFT

### v1

@@ I Introduction
The topic of impossible measurements [sorkin1993impossible] has received renewed attention over the past few years. A multitude of approaches have been developed to address the impossible measurement scenarios (see [papageorgiou2024eliminating] and references therein). The clear moral that has emerged is that the set of quantum mechanical operations, or channels, typically allowed in quantum information is far too large when brought to bear on relativistic theories. In the other direction, when one formulates local measurement theory for relativistic quantum field theory (QFT), the resulting channels do not necessarily include the most familiar ones from quantum information.
A straightforward implication is that the quantum informational machinery of states and operations cannot be straightforwardly imported into QFT. In fact, the established empirical success of QFT relies on the asymptotic scattering paradigm, which is not suitable for describing local operations, i.e. operations over bounded regions of spacetime (for historical context see [Fraser2023]). Sorkin’s examples of impossible operations [sorkin1993impossible], and recent refinements [Borsten:2019cpc], showed that operations built out of local operators (i.e. local Kraus operators) generally allow superluminal signalling. Specifically, this includes ideal measurements of many important observables [Benincasa_2014, albertini2023ideal, PhysRevD.105.025003], leading to a number of papers suggesting non-ideal measurement channels, both abstract [PhysRevD.105.025003, Gisin2024towardsmeasurement, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and concrete i.e. derived from some purified model [mandrysch2024quantumfieldmeasurementsfewsterverch]. The former use a number of no-signalling conditions on the abstract channels or their Kraus operators, while the latter rely on properties of the purified dynamics.
From the opposite direction, a rigorous framework for quantum measurement theory in algebraic QFT has been proposed by Fewster and Verch (FV framework)[fewster2020quantum, PhysRevD.103.025017, Fewster2023], based on purified dynamics. Together, these results point towards a general theory of measurement and operations in QFT. However, there is still no general set of operational criteria, in terms of e.g. Kraus operators and or their purifications, that characterises which measurement maps and operations are physically implementable. A fundamental step towards this would be a “causal Stinespring theorem”, an idea on which we comment at the end of this paper.
The main results of the paper are as follows. We adapt the local $S$-matrix formalism due to Bogoliubov [pittphilsci24053] to the setting of Sorkin’s impossible operations. We demonstrate which of the factorisation conditions are necessary or sufficient for blocking impossible operations, faster than light signalling and retrocausality. In particular, we demonstrate the sufficiency of the strongest factorisation condition, continuous additivity of the local $S$-matrix
@@
Finally, we use an exactly solvable concrete measurement model that can seen as a relativistic von Neumann measurement model, where a local pointer variable is linearly coupled to a smeared scalar field operator, as $\phi(f)\otimes P$ [PhysRevD.109.065024]. We find that the $S$-matrices and Kraus operators obey the conditions (1) and (2), and so this forms a concrete example of the link between the two representations of channels and no-signalling conditions. Further, by solving the local $S$-matrix model we obtain an explicit class of measurement channels on the field whose sharpness is fundamentally limited by the retarded field propagation within the support of $f$, thereby imposing an intrinsic bound on the accuracy of local field measurements. These channels coincide with the abstract continuous measurement channels studied in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], once this new sharpness constraint is taken into account.
@@
The paper is structured as follows: In section II we briefly introduce standard measurement theory and a minimal (algebraic) QFT set-up. We review an abstract notion of operation in the QFT set-up, and briefly present the challenges posed by the impossible measurements scenario, using a simple example of impossible unitary operations introduced in [PhysRevD.105.025003]. In section III we introduce the local $S$-matrix formalism, to discuss the properties of relativistically valid unitary operations in QFT that can (in principle) represent local (not asymptotic) interactions. We take the local $S$-matrix to represent the coupling of the field to a probe system [hellwigkraus1970b] and we examine the operational consequences of a hierarchy of factorisation properties of the $S$-matrix. In section IV we explicitly introduce field operations induced by probes, and the induced Kraus operators in the field’s Hilbert space that are fixed by the local $S$-matrix and the characteristics of the probe system. We consider a simple model for measurements of $\phi(f)$, and consider the limitations to the accuracy of field measurements that come from the causal (retarded) field propagation within the support of $f$. We show how this model relates to the families of Kraus operators for field measurements introduced in Albertini and Jubb [albertini2023ideal] and Oeckl [oeckl2024spectraldecompositionfieldoperators]. Specifically, the result that the sharpness of a measurement of $\phi(f)$ is bounded from below by $\Delta_{\textsc{r}}(f,f)^{1/2}$, i.e. the square-root of the retarded Green’s function smeared against $f$, does not appear in previously suggested measurement maps. In section V we derive no-signalling properties of the induced Kraus operators that can ensure no superluminal signalling and no retrocausality. We discuss the necessity/sufficiency of these conditions and for what type of models they are expected to hold. Finally, we discuss the implications of this model for measurements of more general QFT observables and how this analysis provides hints for relativistic dilation theorems in QFT.
@@ II Local operations in QFT
Perhaps surprisingly, local measurement theory for QFT is still being developed today. This is because the standard predictions and experiments of QFT take the form of asymptotic scattering amplitudes, yet this is not sufficient for modelling local experiments in a relativistic setting. The main obstacle in extracting local probabilistic predictions from QFT theory is that standard quantum mechanical prescriptions for ideal, or projective, measurements and unitary kicks cannot be straightforwardly applied. These operations, when extended to QFT, can lead to signalling across spacelike separated regions. The problem of impossible operations, which was first introduced by Sorkin [sorkin1993impossible], was meant to highlight the friction between no-signalling and the quantum operations that are typically assumed in quantum mechanics. For examples see [Benincasa_2014] and more recently [PhysRevD.105.025003, albertini2023ideal].
@@ II.1 Standard measurement theory
Quantum mechanics has a standard measurement theory, see [busch_book_2016, busch_standard_1996]. Despite ongoing ambiguity regarding what precisely qualifies certain interactions as measurements – a central aspect of the quantum measurement problem – there exist standard prescriptions for extracting probabilistic predictions for any measurement or general operation defined within the theory. In this subsection we recall the basic ideas of measurement theory, in preparation for application to the QFT setup.
@@ II.2 Set-up
Below we introduce the set-up and we review an example from [PhysRevD.105.025003] that is most relevant for the measurement model that we will consider in section IV.2. We focus on the real scalar field. Finally, we introduce the notion of a local operation, which is key for our understanding of Sorkin’s impossible operations.
@@
Taking the Weyl operators with $\operatorname{supp}{f}\subset$ R as generators we can construct a Weyl algebra for each region R, denoted $\mathcal{W}(\text{R})$. Finally, we can take the double commutant $\mathfrak{A}(\text{R}):=\mathcal{W}(\text{R})^{\prime\prime}$—where the commutant of $S\subseteq\mathcal{B}(\mathcal{H})$ is the set of all operators which commute with all elements of SS–––which associates a concrete von Neumann algebra to every region R. We consider an operator $A$ to be localised [^1] AA is localised to R if $A$ is affiliated to $\mathfrak{A}(\text{R})$. to R if $A$ commutes with all operators spacelike to R, i.e. $A,\mathfrak{A}(\text{R}^{\prime})]=0$ where $\text{R}^{\prime}$ is the set of all points spacelike to R. An example is $\phi(f)$, which is local to $\operatorname{supp}{f}\subset\text{R}$, despite being unbounded and thus not in $\mathfrak{A}(\text{R})$. It follows that any bounded function of $\phi(f)$, defined by the spectral calculus, is in $\mathfrak{A}(\text{R})$. For more details on this construction, see [fewster2019algebraic]. Using the properties of the commutator, it follows, from strong commutativity of the fields [Borchers:1992rq], that spacelike algebras commute,
@@
where regions $R$ and $R^{\prime}$ are space-like separated. More generally, in axiomatic approaches to QFT (e.g. algebraic QFT [haag_kastler_1964]) one postulates the axiom of microcausality, namely that spacelike separated algebras of observables commute. This is a kinematical ‘separability’ assumption that, roughly, splits the field degrees of freedom between causally disconnected parts [earman2014relativistic]. Note that microcausality alone does not have an a priori operational consequence [Ruetsche2021, Calderon2022, soulas2025proof] e.g. in terms of signalling. Once a notion of operation is introduced, microcausality alone is generally not sufficient for blocking superluminal signalling (see also [redei_how_2010]). Sorkin showed that [sorkin1993impossible], despite the microcausality assumption, even the most basic local operations in QFT, such as unitary kicks and ideal measurements – when followed by the Lüders post-measurement state update for non-selective measurements – can lead to superluminal signaling between spacelike separated regions (see also [Borsten:2019cpc, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]).
@@ II.3 Impossible operations
The set-up of the Sorkin problem is the following: consider three regions $\text{A},\text{B},\text{C}$ (figure 1) singled out by the supports of the smearing functions $h,f,g$ respectively. The regions are such that with respect to suitable partial ordering $\text{A}\prec\text{B}\prec\text{C}$ (since B is partially in the causal future of A and C is partially in the causal future of B and the partial order is not transitive), but A and C are spacelike separated. Two operations $\Phi_{\textsc{{a,b}}}$ are applied over regions A and B. For region A to not signal to region C, the expectation values of observables in C must not depend on the choice $\lambda$ of which operation $\Phi^{\lambda}_{\textsc{a}}$ is implemented in A. Namely, the condition for no impossible operations (which has previously been considered in [Borsten:2019cpc, deramon2021relativistic, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]) is simply the following: let $\rho$ be the global state of the field and $C$ an observable over region C, then it should hold that
@@ II.3.1 Example of signalling and non-signalling unitaries
The most basic example of failure of this condition using unitary operators in QFT was given in [PhysRevD.105.025003] and uses the real scalar field construction given above. Letting
@@
The failure of the no-signalling condition (21) in Eq (23) can be traced back to the fact that $\Phi_{\textsc{b}}(C)$ will not commute with $A$ in general, even though $[A,C]=0$ by microcausality of operators in AQFT [Borsten:2019cpc], while commutation does hold for Eq (25).
@@ III Local $S$-matrix formalism
Bogoliubov’s local $S$-matrix  [^4] The first attempt to formulate relativistic quantum theory in terms of the $S$-matrix is due to Heisenberg, who introduced a ‘global’ $S$-matrix satisfying unitarity and Lorentz invariance. The local $S$-matrix approach was later introduced by Stueckelberg and Bogoliubov, reintroducing local degrees of freedom and demanding additional causality conditions [pittphilsci24053]. formalism [bogoliubov1960introduction] is a local and covariant approach to describing interactions in QFT, and so is a natural place for us to start. It plays a major role in perturbative algebraic QFT (see [rejzner2019localitycausalityperturbativeaqft] and references within). A local $S$-matrix will be denoted as $S[f]$ where $f$ a compactly supported smearing function. In some cases, when the adiabatic limit can be taken $f(\mathsf{x})\rightarrow f_{\text{const}}$, we end up with the usual global (asymptotic) $S$-matrix familiar from textbook QFT, where coupling terms are spacetime independent constants e.g. $\mathcal{L}_{I}(\mathsf{x})=-f_{\text{const}}\phi(\mathsf{x})^{4}$, and which map between the in and out Hilbert spaces. While the local $S$-matrix formalism was originally conceived, and still used in perturbative AQFT (pAQFT), to define scattering interactions in cases where the asymptotic $S$-matrix might not be well defined, it also provides a useful model for more general localised interactions, which may include scenarios beyond scattering. Further, as we will see, it comes endowed with a hierarchy of causal conditions which are very well suited for studying causality in finite time processes. For these reasons, we will be concerned only with the local case, where the support of $f$ is bounded.
@@
One natural way to impose no-signalling, involves zooming into the support of $f$ and make sure that the operation can be decomposed such that future evolution cannot affect past evolution (retrocausality).  [^5] This condition can be imposed on a Dyson expansion of an $S$-matrix by assuming the unsmeared Lagrangian is microcausal. We call continuous additivity [bogoliubov1960introduction] the following property: given a hypersurface $\Sigma$ that intersects the region $\text{supp}f$ (see figure 3) we consider the decomposition $f=f_{+}+f_{-}$ where $\text{supp}f_{\pm}=\text{supp}f\cap J^{\pm}(\Sigma)$. Then
@@
This property allows us to decompose the dynamical evolution in past and future evolution with respect to any given foliation.
@@ III.1 Properties of valid local $S$-matrices
A few comments are in order regarding the meaning of Eq (31), since this property is central in our analysis of (retro)causality [^7] More abstractly, in local operational frameworks (see [oeckl_local_2019]) one demands a similar compositionality property of null probes (that represent dynamics) that share a boundary. . For our purposes, we demand (31) inspired by the causal perturbation theory program888where the condition is normally introduced in order to help defining the local algebras of a perturbative and causal QFT, see [Ilyin:1978gq, Brunetti_2000, rejzner2019localitycausalityperturbativeaqft] and a recent project for non-perturbative AQFT [Buchholz_2020]. In these works, it is analysed in detail how restrictive this assumption is for realistic interactions, which of course goes beyond our purposes., as well as weaker factorisation conditions (see section VII.1) to analyse their operationally meaningful consequences, in the context of signalling, in a simple measurement scenario involving a probe locally coupled to a field. A final comment about (31) is that it is key that we are cutting $f$ on a Cauchy slice, i.e. a spacelike hypersurface, rather than a lightlike hypersurface such as one that coincides with A or C’s lightcones, as factorising the local $S$-matrix on the lightcone would introduce divergences (which can be seen using microlocal analysis [Brunetti_2000]).

### v2

@@ I Introduction
The topic of impossible measurements [47] has received renewed attention over the past few years. A multitude of approaches have been developed to address the impossible measurement scenarios (see [38] and references therein). The clear moral that has emerged is that the set of quantum mechanical operations, or channels, typically allowed in quantum information is far too large when brought to bear on relativistic theories. In the other direction, when one formulates local measurement theory for relativistic quantum field theory (QFT), the resulting channels do not necessarily include the most familiar ones from quantum information.
A straightforward implication is that the quantum informational machinery of states and operations cannot be straightforwardly imported into QFT. In fact, the established empirical success of QFT relies on the asymptotic scattering paradigm, which is not suitable for describing local operations, i.e. operations over bounded regions of spacetime (for historical context see [19]). Sorkin’s examples of impossible operations [47], and recent refinements [6], showed that operations built out of local operators (i.e. local Kraus operators) generally allow superluminal signalling. Specifically, this includes ideal measurements of many important observables [2, 1, 28], leading to a number of papers suggesting non-ideal measurement channels, both abstract [28, 22, 1, 35] and concrete i.e. derived from some purified model [30]. The former use a number of no-signalling conditions on the abstract channels or their Kraus operators, while the latter rely on properties of the purified dynamics.
From the opposite direction, a rigorous framework for quantum measurement theory in algebraic QFT has been proposed by Fewster and Verch (FV framework)[16, 7, 17], based on purified dynamics. Together, these results point towards a general theory of measurement and operations in QFT. However, there is still no general set of operational criteria, in terms of e.g. Kraus operators and or their purifications, that characterises which measurement maps and operations are physically implementable. A fundamental step towards this would be a “causal Stinespring theorem”, an idea on which we comment at the end of this paper.
The main results of the paper are as follows. We adapt the local $S$-matrix formalism due to Bogoliubov [4, 3] to the setting of Sorkin’s impossible operations. We demonstrate which of the factorisation conditions are necessary or sufficient for blocking impossible operations, faster than light signalling and retrocausality. In particular, we demonstrate the sufficiency of the strongest factorisation condition, continuous additivity of the local $S$-matrix
@@
Finally, we use an exactly solvable concrete measurement model that can seen as a relativistic von Neumann measurement model, where a local pointer variable is linearly coupled to a smeared scalar field operator, as $\phi(f)\otimes P$ [37]. We find that the $S$-matrices and Kraus operators obey the conditions (1) and (2), and so this forms a concrete example of the link between the two representations of channels and no-signalling conditions. Further, by solving the local $S$-matrix model we obtain an explicit class of measurement channels on the field whose sharpness is fundamentally limited by the retarded field propagation within the support of $f$, thereby imposing an intrinsic bound on the accuracy of local field measurements. These channels coincide with the abstract continuous measurement channels studied in [28, 1, 35], once this new sharpness constraint is taken into account.
@@
The paper is structured as follows: In section II we briefly introduce standard measurement theory and a minimal (algebraic) QFT set-up. We review an abstract notion of operation in the QFT set-up, and briefly present the challenges posed by the impossible measurements scenario, using a simple example of impossible unitary operations introduced in [28]. In section III we introduce the local $S$-matrix formalism, to discuss the properties of relativistically valid unitary operations in QFT that can (in principle) represent local (not asymptotic) interactions. We take the local $S$-matrix to represent the coupling of the field to a probe system [26] and we examine the operational consequences of a hierarchy of factorisation properties of the $S$-matrix. In section IV we explicitly introduce field operations induced by probes, and the induced Kraus operators in the field’s Hilbert space that are fixed by the local $S$-matrix and the characteristics of the probe system. We consider a simple model for measurements of $\phi(f)$, and consider the limitations to the accuracy of field measurements that come from the causal (retarded) field propagation within the support of $f$. We show how this model relates to the families of Kraus operators for field measurements introduced in Albertini and Jubb [1] and Oeckl [35]. Specifically, the result that the sharpness of a measurement of $\phi(f)$ is bounded from below by $\Delta_{\textsc{r}}(f,f)^{1/2}$, i.e. the square-root of the retarded Green’s function smeared against $f$, does not appear in previously suggested measurement maps. In section V we derive no-signalling properties of the induced Kraus operators that can ensure no superluminal signalling and no retrocausality. We discuss the necessity/sufficiency of these conditions and for what type of models they are expected to hold. Finally, we discuss the implications of this model for measurements of more general QFT observables and how this analysis provides hints for relativistic dilation theorems in QFT.
@@ II Local operations in QFT
Perhaps surprisingly, local measurement theory for QFT is still being developed today. This is because the standard predictions and experiments of QFT take the form of asymptotic scattering amplitudes, yet this is not sufficient for modelling local experiments in a relativistic setting. The main obstacle in extracting local probabilistic predictions from QFT theory is that standard quantum mechanical prescriptions for ideal, or projective, measurements and unitary kicks cannot be straightforwardly applied. These operations, when extended to QFT, can lead to signalling across spacelike separated regions. The problem of impossible operations, which was first introduced by Sorkin [47], was meant to highlight the friction between no-signalling and the quantum operations that are typically assumed in quantum mechanics. For examples see [2] and more recently [28, 1].
@@ II.1 Standard measurement theory
Quantum mechanics has a standard measurement theory, see [10, 11]. Despite ongoing ambiguity regarding what precisely qualifies certain interactions as measurements – a central aspect of the quantum measurement problem – there exist standard prescriptions for extracting probabilistic predictions for any measurement or general operation defined within the theory. In this subsection we recall the basic ideas of measurement theory, in preparation for application to the QFT setup.
@@ II.2 Set-up
Below we introduce the set-up and we review an example from [28] that is most relevant for the measurement model that we will consider in section IV.2. We focus on the real scalar field. Finally, we introduce the notion of a local operation, which is key for our understanding of Sorkin’s impossible operations.
@@
Taking the Weyl operators with $\operatorname{supp}{f}\subset$ R as generators we can construct a Weyl algebra for each region R, denoted $\mathcal{W}(\text{R})$. Finally, we can take the double commutant $\mathfrak{A}(\text{R}):=\mathcal{W}(\text{R})^{\prime\prime}$—where the commutant of $S\subseteq\mathcal{B}(\mathcal{H})$ is the set of all operators which commute with all elements of SS–––which associates a concrete von Neumann algebra to every region R. We consider an operator $A$ to be localised [^1] AA is localised to R if $A$ is affiliated to $\mathfrak{A}(\text{R})$. to R if $A$ commutes with all operators spacelike to R, i.e. $A,\mathfrak{A}(\text{R}^{\prime})]=0$ where $\text{R}^{\prime}$ is the set of all points spacelike to R. An example is $\phi(f)$, which is local to $\operatorname{supp}{f}\subset\text{R}$, despite being unbounded and thus not in $\mathfrak{A}(\text{R})$. It follows that any bounded function of $\phi(f)$, defined by the spectral calculus, is in $\mathfrak{A}(\text{R})$. For more details on this construction, see [18]. Using the properties of the commutator, it follows, from strong commutativity of the fields [5], that spacelike algebras commute,
@@
where regions $R$ and $R^{\prime}$ are space-like separated. More generally, in axiomatic approaches to QFT (e.g. algebraic QFT [23]) one postulates the axiom of microcausality, namely that spacelike separated algebras of observables commute. This is a kinematical ‘separability’ assumption that, roughly, splits the field degrees of freedom between causally disconnected parts [15]. Note that microcausality alone does not have an a priori operational consequence [44, 12, 48] e.g. in terms of signalling. Once a notion of operation is introduced, microcausality alone is generally not sufficient for blocking superluminal signalling (see also [41]). Sorkin showed that [47], despite the microcausality assumption, even the most basic local operations in QFT, such as unitary kicks and ideal measurements – when followed by the Lüders post-measurement state update for non-selective measurements – can lead to superluminal signaling between spacelike separated regions (see also [6, 28, 35]).
@@ II.3 Impossible operations
The set-up of the Sorkin problem is the following: consider three regions $\text{A},\text{B},\text{C}$ (figure 1) singled out by the supports of the smearing functions $h,f,g$ respectively. The regions are such that with respect to suitable partial ordering $\text{A}\prec\text{B}\prec\text{C}$ (since B is partially in the causal future of A and C is partially in the causal future of B and the partial order is not transitive), but A and C are spacelike separated. Two operations $\Phi_{\textsc{{a,b}}}$ are applied over regions A and B. For region A to not signal to region C, the expectation values of observables in C must not depend on the choice $\lambda$ of which operation $\Phi^{\lambda}_{\textsc{a}}$ is implemented in A. Namely, the condition for no impossible operations (which has previously been considered in [6, 14, 28, 35]) is simply the following: let $\rho$ be the global state of the field and $C$ an observable over region C, then it should hold that
@@ II.3.1 Example of signalling and non-signalling unitaries
The most basic example of failure of this condition using unitary operators in QFT was given in [28] and uses the real scalar field construction given above. Letting
@@
The failure of the no-signalling condition (21) in Eq (23) can be traced back to the fact that $\Phi_{\textsc{b}}(C)$ will not commute with $A$ in general, even though $[A,C]=0$ by microcausality of operators in AQFT [6], while commutation does hold for Eq (25).
@@ III Local $S$-matrix formalism
Bogoliubov’s local $S$-matrix  [^4] The first attempt to formulate relativistic quantum theory in terms of the $S$-matrix is due to Heisenberg, who introduced a ‘global’ $S$-matrix satisfying unitarity and Lorentz invariance. The local $S$-matrix approach was later introduced by Stueckelberg and Bogoliubov, reintroducing local degrees of freedom and demanding additional causality conditions [3]. formalism [4] is a local and covariant approach to describing interactions in QFT, and so is a natural place for us to start. It plays a major role in perturbative algebraic QFT (see [42] and references within). A local $S$-matrix will be denoted as $S[f]$ where $f$ a compactly supported smearing function. In some cases, when the adiabatic limit can be taken $f(\mathsf{x})\rightarrow f_{\text{const}}$, we end up with the usual global (asymptotic) $S$-matrix familiar from textbook QFT, where coupling terms are spacetime independent constants e.g. $\mathcal{L}_{I}(\mathsf{x})=-f_{\text{const}}\phi(\mathsf{x})^{4}$, and which map between the in and out Hilbert spaces. While the local $S$-matrix formalism was originally conceived, and still used in perturbative AQFT (pAQFT), to define scattering interactions in cases where the asymptotic $S$-matrix might not be well defined (for historical context see [20, 3]), it also provides a useful model for more general localised interactions, which may include scenarios beyond scattering. Further, as we will see, it comes endowed with a hierarchy of causal conditions which are very well suited for studying causality in finite time processes. For these reasons, we will be concerned only with the local case, where the support of $f$ is bounded.
@@
One natural way to impose no-signalling, involves zooming into the support of $f$ and make sure that the operation can be decomposed such that future evolution cannot affect past evolution (retrocausality).  [^5] This condition can be imposed on a Dyson expansion of an $S$-matrix by assuming the unsmeared Lagrangian is microcausal. We call continuous additivity [4] the following property: given a hypersurface $\Sigma$ that intersects the region $\text{supp}f$ (see figure 3) we consider the decomposition $f=f_{+}+f_{-}$ where $\text{supp}f_{\pm}=\text{supp}f\cap J^{\pm}(\Sigma)$. Then
@@
This property allows us to decompose the dynamical evolution in past and future evolution with respect to any given foliation [^6] Note that by decomposing $f$ in this way, $f_{\pm}$ will generally not be smooth. Any mathematical problems that may arise from this (e.g. surface divergences, see [3, 20]) do not arise in the more complex Hammerstein factorisation, where all functions can be kept $C^{\infty}$ smooth. .
@@ III.1 Properties of valid local $S$-matrices
A few comments are in order regarding the meaning of Eq (31), since this property is central in our analysis of (retro)causality [^8] More abstractly, in local operational frameworks (see [33]) one demands a similar compositionality property of null probes (that represent dynamics) that share a boundary. . For our purposes, we demand (31) inspired by the causal perturbation theory program999where the condition is normally introduced in order to help defining the local algebras of a perturbative and causal QFT, see [27, 8, 42] and a recent project for non-perturbative AQFT [9]. In these works, it is analysed in detail how restrictive this assumption is for realistic interactions, which of course goes beyond our purposes., as well as weaker factorisation conditions (see section VII.1) to analyse their operationally meaningful consequences, in the context of signalling, in a simple measurement scenario involving a probe locally coupled to a field. A final comment about (31) is that it is key that we are cutting $f$ on a Cauchy slice, i.e. a spacelike hypersurface, rather than a lightlike hypersurface such as one that coincides with A or C’s lightcones, as factorising the local $S$-matrix on the lightcone would introduce divergences (which can be seen using microlocal analysis [8]).